A solid sphere of radius 5 cm floats in water. If a maximum load of 0.1 kg can be put on it without wetting the load, find the specific gravity of the material of the sphere.

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol. W1 + W2 = U. ⇒mg + V * ps * g = V * pw * g (where ps = density of sphere in gm/cc) ⇒1 – ps = 0.19 ⇒ps = 1 – (0.19) = 0.8 gm/cc So, specific gravity of the material is 0.8.